OpenStudy (anonymous):

What is the geometric mean of 3 and 21?

OpenStudy (turingtest):

\[\sqrt{3*21}\]

OpenStudy (anonymous):

with the arithmetic mean we add and then divide by two .... with the geometric mean we multiply and take the square root..

OpenStudy (turingtest):

In general the geometric mean of n terms is given by\[\overline{a}=\sqrt[n]{a_1\times a_2\times a_3\times...\times a_{n-1}\times a_n}\]where\[a_i\]is the i-th term in the sequence.